Three fixed point charges are at the vertices of a
Answer #1


r1 = r2 = sqrt(34.7^2+9.3^2) = 35.9 cm


theta1 = theta2 = tan^-1(9.3/34.7) = 15 degrees


r3 = (34.7*tan30) - 9.3 = 10.6 cm

E1x = k*q1*costheta1/r1^2 = (9*10^9*5*10^-6*cos15)/0.359^2 = 337261.98 N/C

E1y = k*q1*sintheta1/r1^2 = (9*10^9*5*10^-6*sin15)/0.359^2 = 90369.08 N/C


E2x = -k*q2*costheta1/r1^2 = -(9*10^9*5*10^-6*cos15)/0.359^2 = -337261.98 N/C


E2y = k*q2*sintheta1/r1^2 = (9*10^9*5*10^-6*sin15)/0.359^2 = 90369.08 N/C

E3y = -k*q3/r3^2 = -(9*10^9*5*10^-6)/0.106^2 = -4004983.98 N/C


Ex = E1x + E2x +E3x = 0

Ey = E1y + E2y + E3y

Ey = -3824245.82 N/C

net force E = -3824245.82 N/C


net force F = E*Q = -3824245.82*20*10^-6 = -76.5 N

magnitude = 76.5 N

direction = down wards